OpenStudy (x__hazel__x):

What is 50 cents + (25 cents × 4) + (10 cents × 7) + (5 cents × 58)? I would also appreciate it if you could answer using PEMDAS, but if you cant, that is fine :3

OpenStudy (anonymous):

First do the items in each of the parenthesis. What do you get?

OpenStudy (mervatm):

25c*4=1.00 10c*7=70c 5c*58=2.90 1.00+2.90=3.90 3.90+70=4.60 4.60+50c=5.10 =$5.10

OpenStudy (x__hazel__x):

Thank you!